A +11 nC charge is located at (x,y)=(0cm,12cm) and a -3.0 nC charge is located at (x,y)=(7.0cm,0cm).

Where would a -14 nC charge need to be located in order that the electric field at the origin be zero?
Answer #1

.11nC 12cm 3nC /cm

Electric field ,E=kQ/r^{2}

E, = k * llncy(0.12771 )

El-9 * 10% * 1.1 * 10-8/0.122

E, = 6875N/C

{color{Red} vec{E_{1}}=-6875hat{j} N/C}

===================

E-2-k * 3nC/ (0.07 m)2

Ег 9 * 109 * 3 * 10-9/0.072

E, = 5510 .2Ус

E2 = 5510 .2ì N/C

===================

vec{E_{3}}=E_{3x}hat{i}+E_{3y}hat{j}

Given vec{E_{1}}+vec{E_{2}}+vec{E_{3}}=0

-6875hat{j}+5510.2hat{i}+E_{3x}hat{i}+E_{3y}hat{j}=0

(5510.2+E_{3x})hat{i}+(-6875+E_{3y})hat{j}=0

(5510.2+E_{3x})=0

E_{3x}=-5510.2

kq_{3}/x^{2}=-5510.2

9*10^{9}*1.4*10^{-8}/x^{2}=-5510.2

9*10^{9}*1.4*10^{-8}/5510.2=x^{2}

Ignore negative sign .it just shows x is negative

x=0.1512m=15.12cm

ANSWER: {color{Red} x=-15.12cm}

===================

Similarly

(5510.2+E_{3x})hat{i}+(-6875+E_{3y})hat{j}=0

-6875+E_{3y}=0

6875=E_{3y}

6875=kq_{3}/y^{2}

6875-9 * 109 * 1.4 * 10-8/y2

y^{2}=9*10^{9}*1.4*10^{-8}/6875

y=0.1354m=13.54cm

ANSWER: {color{Red} y=13.54cm}

====================

ANSWER: {color{Red} (-15.12cm,13.54cm)}

================================